Find how much should be deposited now at 8% interest, compounded semiannually, to yield an annuity payment of $700 at the BEGINNING of each 6 months, for 2 years
Answer #1

Semi-annual interest = 8%/2 = 4%

Number of periods = 2*2 = 4

Amount to be deposited today is equal to present value of withdrawals

= 700 + 700*PVAF(4%, 3 periods)

= 700 + 700*2.7751

= $2,642.57